# Prediction of linear fractional stable motions using codifference

### v1

@@ 3.2 From the decomposition to the forecast
We assume we observe the process at $d-1$ discrete times: $X_{1},...,X_{d-1}$. We want to forecast $\mathcal{T}X_{d}$. As exposed in Section 3.1, it can be extended to cases where all the times are translated by $t$.
@@ 3.2.1 Conditional expectation: $\alpha\in(1,2]$
where $\mathbf{Z}_{0,d-1}=(Z_{0},...,Z_{d-1})^{\prime}$ is a jointly $S\alpha S$ independent vector of unitary scale parameter and the matrix $\mathbf{A}_{0,d-1}\in\mathbb{R}^{d\times d}$, of element $[\mathbf{A}_{0,d-1}]_{ij}=a_{1,i-1,j-1}$, is defined in accordance with the details given in Section 3.1.
Inspired by the Gaussian case [53], one can be tempted to build a forecast of $X_{d}$ by considering its conditional expectation. For $S\alpha S$ variables, the conditional expectation is well defined for $\alpha>1$. When $\alpha\leq 1$, it can also be well defined under some restrictive conditions [60, Chapter 5]. In our case, since $\mathcal{T}X_{d}$ is obtained by adding an independent increment to a linear combination of $\mathcal{T}X_{1}$, …, $\mathcal{T}X_{d-1}$, the conditional expectation $\mathbb{E}[\mathcal{T}X_{d}|\mathbf{X}_{1,d-1}]$ only exists for $\alpha>1$. In this case, by linearity of the expectation, equation (7) and $\mathbf{X}_{1,d-1}=\mathbf{A}_{0,d-2}\mathbf{Z}_{0,d-2}$ give the following forecast:
@@
We now summarize the steps of the cascade algorithm, which leads to the forecast $\widehat{\mathcal{T}X}_{d}$. For $i$ successively equal to 0, …, $d-2$, we do the following:
@@ 2. Step 2:
At the end, we also compute step 1 for $i=d-1$, so that we can calculate $\widehat{\mathcal{T}X}_{d}$ using equation (8).
In step 2, we use $X_{i+1}$, which is observed. When working conditionally on past observations $X_{1}$, …, $X_{d}$, we assume that $\mathcal{T}X_{i}=X_{i}$ for $i\leq d-1$. We may only have a divergence between $\mathcal{T}X_{d}$ and $X_{d}$, which are both unobserved at this date. In other words, we forecast $\mathcal{T}X_{d}$ based on a model that might be slightly different from the one of $X_{d}$,111 They only have the same codifference, not necessarily the same spectral measure. but using exactly the same conditioning on past observations.
@@ 3.3 Quality of the forecast
[…]
From this, we easily get the $p$-th absolute moment of $X\sim\mathcal{N}(0,1)$, when $p>-1$:
[…]
@@ 4 Simulation study
The hit ratio largely below $50\%$, which we obtain for very low values of $\alpha$, can be seen as a paradox, but we are able to explain it. For an fBm with $H<1/2$, non-overlapping increments are negatively correlated. Considering three consecutive increments, if the first one is positive, the second one, as negatively correlated to the first one, is more likely to be negative. The third one is thus negatively correlated to a positive and to a negative increment, but the correlation decreases rapidly in this framework where there is no long-range dependence. Therefore, the third increment will more likely be positive. In the case of an LFSM with $\alpha$ below $1/H$, the alternation of positive and negative increments is also very likely. But when, in addition, $\alpha$ is very small, the frequent occurrence of very large increments can disrupt the mechanism described above. Indeed, when one observe a very large increment, say a positive one, the next increment will more likely be a large, but not as large, negative number. Then, the third increment should be positive due to its negative codifference with the second one. But the first increment is so large that the negative codifference between the first and third increments will dominate the codifference between the second and third ones. Consequently, the third increment will more likely be negative. So the second and third increments give the illusion of being positively dependent, what we can explain, in causality terms, by the presence of an overwhelming confounder, namely the first and large increment. At the macroscopic scale, after a large positive increment, subsequent increments will be negative and their magnitude will gradually decrease. When considering the process itself instead of the increments, one thus observes after each large jump a kind of local trend or progressive recovery. This phenomenon is visible in the simulation displayed in the bottom left graph of Figure 1.
In the literature, $\alpha<1/H$ is associated to an absence of long-rang dependence [60, Section 7.4]. But the explanation above shows that large increments, which occur for small values of $\alpha$, may disrupt this interpretation. This justifies why a forecast based on a limited number of past observations, thus neglecting some past large increments, performs poorly when $\alpha$ is very small.
@@ 5 Empirical application
In this section, we investigate the performance of the forecasting method based on LFSM when applied to real data. We focus on financial data, namely time series of realized volatilities and time series of FX-rates sampled every minute.
@@ 5.1 Rough volatility
We propose here to replace the Gaussian distribution of the fBm by an $\alpha$-stable one and to apply the forecasting method of the LFSM exposed in Section 3.2 to time series of volatilities. The data used in our analysis are the same as those used in a previous work on forecasting volatilities with an fBm [28], namely daily realized volatilities computed with a five-minute discretization of prices, imported from the formerly available Oxford-Man Institute of Quantitative Finance Realized Library. We focus on the realized volatility of eight stock indices: the AEX index, the CAC 40 index, the FTSE 100 index, the Nasdaq 100 index (IXIC), the Nikkei 225 index (N225), the Oslo Exchange All-share index (OSEAX), the Madrid General index (SMSI), and the S&P 500 index. The series starts on January 2000, except N225, which starts in February 2000, OSEAX in September 2001 and SMSI in July 2005. The end date of our sample is on the 12th April 2021. The purpose of the study is to forecast the next daily variation of volatility. For each series, and each day $t$, we estimate the parameters of an LFSM in the two-year window finishing at $t$. Next, using these parameters and the method developed in Section 3.2, we forecast the volatility of day $t+1$.
@@
Figure 6 displays the trajectory of the realized volatility of the CAC 40 index, along with the estimates of the parameters $\alpha$ (between 1.65 and 2) and $H$ (lower than $1/2$) of an LFSM. We also represent $H-1/\alpha$, that is the memory of the process, which, as one can see in Figure 6, is always negative. It means that the sign of the forecast of the future increment of volatility is simply the opposite of the one of the last volatility increment, if the forecast is based on this sole lagged observation. In other words, when $d=2$ the LFSM and the fBm will lead to the same hit ratios.
@@ 5.2 High-frequency FX rates
Many papers have already studied the relevance of the fBm for modelling FX rates [10, 66], along with forecasts consistent with this model [24, 28]. Figure 8 represents the three time series of FX rates. It clearly presents some jumps, which legitimize using a stable process instead of a Gaussian one. We then estimate the parameters of an LFSM on the log of each time series, using a sliding window of 720 observations, that is 12 hours. As one can see in Figure 8, the value of $\alpha$ is in general between $1.2$ and 2, with a peak below 1 for GBPUSD on the 26th June, $H$ fluctuates around $1/2$, with a negative peak of the estimator for GBPUSD on the 26th June. The value of the memory, $H-1/\alpha$, is more often negative than positive. We remark that $H-1/2$ and $H-1/\alpha$ do not always have the same sign. This indicates that the forecast using the LFSM is not always the same as the one using the fBm, even for $d=2$.

### v2

@@ 3.2 From the decomposition to the forecast
We assume we observe the process at d1d-1 discrete times: $X_{1},...,X_{d-1}$. We want to forecast $\mathcal{T}X_{d}$. As exposed in Section 3.1, it can be extended to cases where all the times are translated by $t$.
@@ 3.2.1 Conditional expectation: $\alpha\in(1,2]$
where $\mathbf{Z}_{0,d-1}=(Z_{0},...,Z_{d-1})^{\prime}$ is a jointly $S\alpha S$ independent vector of unitary scale parameter and the matrix $\mathbf{A}_{0,d-1}\in\mathbb{R}^{d\times d}$, of element $\mathbf{A}_{0,d-1}]_{ij}=a_{1,i-1,j-1}$, is defined in accordance with the details given in Section 3.1.
Inspired by the Gaussian case [54], one can be tempted to build a forecast of $X_{d}$ by considering its conditional expectation. For $S\alpha S$ variables, the conditional expectation is well defined for $\alpha>1$. When $\alpha\leq 1$, it can also be well defined under some restrictive conditions [61, Chapter 5]. In our case, since $\mathcal{T}X_{d}$ is obtained by adding an independent increment to a linear combination of $\mathcal{T}X_{1}$, …, $\mathcal{T}X_{d-1}$, the conditional expectation $\mathbb{E}[\mathcal{T}X_{d}|\mathbf{X}_{1,d-1}]$ only exists for $\alpha>1$. In this case, by linearity of the expectation, equation (8) and $\mathbf{X}_{1,d-1}=\mathbf{A}_{0,d-2}\mathbf{Z}_{0,d-2}$ give the following forecast:
@@
We now summarize the steps of the cascade algorithm, which leads to the forecast $\widehat{\mathcal{T}X}_{d}$. For $i$ successively equal to 0, …, d2d-2, we do the following:
@@ 2. Step 2:
At the end, we also compute step 1 for i=d1i=d-1, so that we can calculate $\widehat{\mathcal{T}X}_{d}$ using equation (9).
In step 2, we use $X_{i+1}$, which is observed. When working conditionally on past observations $X_{1}$, …, $X_{d}$, we assume that $\mathcal{T}X_{i}=X_{i}$ for $i\leq d-1$. We may only have a divergence between $\mathcal{T}X_{d}$ and $X_{d}$, which are both unobserved at this date. In other words, we forecast $\mathcal{T}X_{d}$ based on a model that might be slightly different from the one of $X_{d}$, [^1] They only have the same codifference, not necessarily the same spectral measure. but using exactly the same conditioning on past observations.
@@ 3.3 Quality of the forecast
$\begin{array}[]{ccl}\int_{0}^{+\infty}x^{s-1}\frac{e^{-x^{2}/2}}{\sqrt{2\pi}}dx&=&\int_{0}^{+\infty}(x^{2})^{(s-2)/2}\frac{e^{-x^{2}/2}}{\sqrt{2\pi}}xdx\\ &=&\int_{0}^{+\infty}(2y)^{(s-2)/2}\frac{e^{-y}}{\sqrt{2\pi}}dy\\ &=&\frac{2^{(s-2)/2}}{\sqrt{2\pi}}\Gamma\left(\frac{s}{2}\right).\end{array}$
From this, we easily get the $p$-th absolute moment of $X\sim\mathcal{N}(0,1)$, when p>1p>-1:
$\mathbb{E}\left(|X|^{p}\right)=2\int_{0}^{+\infty}x^{p}\frac{e^{-x^{2}/2}}{\sqrt{2\pi}}dx=\frac{2^{p/2}}{\sqrt{\pi}}\Gamma\left(\frac{p+1}{2}\right).$
@@ 4 Simulation study
The hit ratio largely below $50\%$ that we obtain for very low values of $\alpha$ can be seen as a paradox, but we are able to explain it. For an fBm with $H<1/2$, non-overlapping increments are negatively correlated. Considering three consecutive increments, if the first one is positive, the second one, as negatively correlated to the first one, is more likely to be negative. The third one is thus negatively correlated to a positive and to a negative increment, but the correlation decreases rapidly in this framework where there is no long-range dependence. Therefore, the third increment will more likely be positive. In the case of an LFSM with $\alpha$ below $1/H$, the alternation of positive and negative increments is also very likely. But when, in addition, $\alpha$ is very small, the frequent occurrence of very large increments can disrupt the mechanism described above. Indeed, when one observes a very large increment, say a positive one, the next increment will more likely be a large, but not as large, negative number. Then, the third increment should be positive due to its negative codifference with the second one. But the first increment is so large that the negative codifference between the first and third increments will dominate the codifference between the second and third ones. Consequently, the third increment will more likely be negative. So the second and third increments give the illusion of being positively dependent, what we can explain, in causality terms, by the presence of an overwhelming confounder, namely the first and large increment. At the macroscopic scale, after a large positive increment, subsequent increments will be negative and their magnitude will gradually decrease. When considering the process itself instead of the increments, one thus observes after each large jump a kind of local trend or progressive recovery. This phenomenon is visible in the simulation displayed in the bottom left graph of Figure 1.
In the literature, $\alpha<1/H$ is associated to an absence of long-rang dependence [61, Section 7.4]. But the explanation above shows that large increments, which occur for small values of $\alpha$, may disrupt this interpretation. This justifies why a forecast based on a limited number of past observations, thus neglecting some past large increments, performs poorly when $\alpha$ is very small.
@@
The above simulation study is based on an oracular framework, in which we use exactly the same parameters to simulate the dataset and to build the LFSM-based forecast. We now present a second and more realistic experiment, in which we ignore the parameters of the generating process. We thus estimate a model and make a forecast based on this estimate.
The generating process is an LFSM of parameters $(\alpha,H)$ successively equal to $(1.5,0.3)$, $(1.5,0.5)$, $(1.5,0.8)$, $(1.8,0.3)$, $(1.8,0.5)$, and $(1.8,0.8)$. For each pair of parameters, we simulate a single trajectory of size 5,000. Then we estimate dynamically three models, in a rolling window of size 500: an LFSM, an fBm, and an autoregressive (AR) model on the increments of the simulated process with d1d-1 lags. The estimation of the AR model is based on a linear regression. The size of each trajectory and of the estimation sample is close to what is done in the empirical study of real volatility data, Section 5. For a given model, we use each estimate and $d$ lagged observations to forecast the observation right after the end of the estimation window, so that, for each trajectory, we have 4,500 out-of-sample point predictions. The accuracy of each model is assessed through the hit ratio and the mean absolute error (MAE) between the forecast and the true value.
Figure 6 gathers the results for $\alpha=1.5$. The three models display a satisfying hit ratio, above $50\%$. When $H\in\{0.3,0.5\}$, case for which the autocodifference is negative, the fBm is the worst of the three models whereas the LFSM and the AR model are close to each other, at least regarding the hit ratio. When $d$ increases, the LFSM outperforms the AR model, both regarding the hit ratio criterion, from $d\geq 12$, and more strongly regarding the MAE, from $d\geq 5$. The poor MAE of the AR model when $d$ is large comes from the overfitting of this model, since the number of its parameters increases with $d$, whereas both the fBm and the LFSM have the same number of parameters whatever $d$. When $H=0.8$, the memory $H-1/\alpha$ is slightly above zero. Since it is close to zero, the uncertainty in the estimation of the parameters of the LFSM may locally lead to an estimate with a negative autocodifference and thus a less good forecast. For this reason, when $(\alpha,H)=(1.5,0.8)$, the LFSM slightly underperforms the two other models, even though, for large values of $d$, the LFSM is on par with the AR model and has a much lower MAE than the fBm.
When $\alpha=1.8$, we see in Figure 7 the same behaviour as the one described above: overall, the LFSM is close to both the fBm and the AR model and it outperforms these two models when $d$ is large. The slightly less favourable case for LFSM is when $H=0.5$, which is the one for which the memory is very slightly positive, like for $(\alpha,H)=(1.5,0.8)$. It indeed again leads to an estimate with uncertain autocodifference. However, when $d$ increases, the MAE criterion indicates that the LFSM is the best of the three models. The parameter set $(\alpha,H)=(1.8,0.5)$ is also the only one in our experiment leading to a hit ratio sometimes lower than $50\%$ and only for the fBm. Indeed, whatever $\alpha$, if $H=0.5$ the estimated fBms have a very fluctuating sign of autocovariance and the resulting predictions are highly uncertain.
Figure 6: Hit ratio (left) and MAE (right) for a trajectory of 5,000 observations generated by an LFSM of parameters $\alpha=1.5$ and $H$ equal to 0.3 (top graphs), 0.5 (middle graphs), and 0.8 (bottom graphs). The series is forecast either with an LFSM (black), or with an fBm (dark grey), or with an AR model (light grey), for $d\in\llbracket 2,15\rrbracket$ after estimation of these models in a rolling window of 500 observations.
Figure 7: Hit ratio (left) and MAE (right) for a trajectory of 5,000 observations generated by an LFSM of parameters $\alpha=1.8$ and $H$ equal to 0.3 (top graphs), 0.5 (middle graphs), and 0.8 (bottom graphs). The series is forecast either with an LFSM (black), or with an fBm (dark grey), or with an AR model (light grey), for $d\in\llbracket 2,15\rrbracket$ after estimation of these models in a rolling window of 500 observations.
@@ 5 Empirical application
In this section, we investigate the performance of the forecasting method based on LFSM when applied to real data. We focus on financial data, namely daily time series of volatilities and time series of FX-rates sampled every minute. We make a special effort to give an empirical evidence of the relevance of this non-Gaussian fractional model in the case of volatility.
@@ 5.1 Rough volatility
We propose here to study the relevance of replacing the Gaussian distribution of the fBm by an $\alpha$-stable one and to apply the forecasting method of the LFSM exposed in Section 3.2 to time series of volatilities. The data used in our analysis are the same as those used in a previous work on forecasting volatilities with an fBm [28], namely daily realized volatilities computed with a five-minute discretization of prices, imported from the formerly available Oxford-Man Institute of Quantitative Finance Realized Library. We focus on the realized volatility of eight stock indices: the AEX index, the CAC 40 index, the FTSE 100 index, the Nasdaq 100 index (IXIC), the Nikkei 225 index (N225), the Oslo Exchange All-share index (OSEAX), the Madrid General index (SMSI), and the S&P 500 index. The series starts on January 2000, except N225, which starts in February 2000, OSEAX in September 2001 and SMSI in July 2005. The end date of our sample is on the 12th April 2021.
The papers that justify the empirical relevance of the fBm for depicting volatilities mainly focus on the fractal properties of the time series and often overlook the justification of the Gaussian distribution. Instead, Table 1 shows that series of the increments of the logarithm of realized volatilities have a significant non-Gaussian behaviour, with a leptokurtic distribution. First, the Jarque-Bera test, which is based on both the skewness and the kurtosis, as well as a Z-test based on the sole kurtosis, both reject with a strong confidence the null hypothesis of Gaussian increments. Second, using the extreme value theory (EVT), we study the speed at which the tails of the density go to zero. For $N$ observations, it is based on those ranked $N-k_{N}$, $N-2k_{N}$, and $N-4k_{N}$, where the conditions $k_{N}\rightarrow\infty$ and $k_{N}/N\rightarrow 0$ ensure the consistency of Pickands estimator $\hat{\xi}_{k_{N}}$ of the EVT parameter. In practice, the estimator is very sensitive to the choice of the free parameter $k_{N}$, so that we define a more robust estimator $\hat{\xi}$ as the median of $\hat{\xi}_{1}$, $\hat{\xi}_{40}$. [^2] In the application, $N$ is around $5,000$, with some variations depending on the series. Simulating $10,000$ times $N$ Gaussian variables, we also get an approximate distribution of these estimators under the null hypothesis of Gaussian increments. Table 1 reveals a distinct behaviour in the right and left tails: although on the left the null hypothesis can only be rejected for SPX, on the right the Fréchet distribution and therefore fat tails are validated for all the series except IXIC. While the observations show some asymmetry, we restrict attention here to the symmetric LFSM, focusing on the impact of heavy tails and serial dependence. Extensions to asymmetric models are left for future work.
series JB statistic kurtosis Z-test statistic EVT right EVT left
AEX $1303^{\star\star\star}$ 5.41 $36.1^{\star\star\star}$ $180^{\star} -0.0$58
CAC 40 $1154^{\star\star\star}$ 5.25 $33.9^{\star\star\star}$ $0.381^{\star\star}$ 0.129
FTSE $416^{\star\star\star}$ 4.35 $20.5^{\star\star\star}$ $181^{\star} -0.0$14
IXIC $233^{\star\star\star}$ 3.97 $14.4^{\star\star\star}$ -0.125 -0.167
N225 $1183^{\star\star\star}$ 5.21 $32.5^{\star\star\star}$ $422^{\star\star\star} 0.0$14
OSEAX $295^{\star\star\star}$ 4.19 $17.0^{\star\star\star}$ $0.179^{\star}$ -0.275
SMSI $1026^{\star\star\star}$ 5.45 $31.7^{\star\star\star}$ $441^{\star\star\star} 0.0$29
SPX $150^{\star\star\star}$ 3.80 $11.8^{\star\star\star}$ $0.478^{\star\star\star}$ $0.233^{\star}$
Table 1: For each series of realized volatility, Jarque-Bera (JB) statistic, kurtosis along with the statistic of the corresponding Z-test, and EVT parameter obtained by Pickands’ estimator, both for right and left tails. The significance at $95\%$ (), $99\%$ (⋆⋆), and $99.9\%$ (⋆⋆⋆) is indicated for the Jarque-Bera test, the Z-test, and the tests on the EVT parameters.
Once the Gaussian distribution is rejected, we can assess the relevance of the $\alpha$-stable alternative by giving some details on the estimation. Indeed, the estimator proposed in Section 2.4 is based on a linear regression of a specific transformation of the empirical characteristic function $\widehat{\Phi}_{X_{.+\tau}-X_{.}}$ of increments, for a reasonable set of parameters $\theta$ and a fixed duration $\tau$. We can see in Figure 8, based on the SPX series, that the linear shape is satisfying, whatever the duration $\tau$, so that one cannot reject the $\alpha$-stable hypothesis. The slope directly provides us with an estimate of $\alpha$, equal to 1.91 in the case the most subject to noise, that is $\tau=1$, and equal to 1.80, 1.82, and 1.77 for $\tau\in\{5,10,22\}$.
Figure 8: $\ln(-\ln(\widehat{\Phi}_{X_{.+\tau}-X_{.}}(\theta)))$ as a function of $\ln|\theta|$, along with its linear regression, for $X$ the full series of the logarithm of the realized volatility on SPX, $\theta\in[0.5,2.5]$, and, from bottom to top, with a duration of increment $\theta$ equal (in days) to 1, 5, 10, and $2$.
The second purpose of this empirical study is to forecast the next daily variation of volatility. For each series, and each day $t$, we estimate the parameters of an LFSM in the two-year window finishing at $t$. Next, using these parameters and the method developed in Section 3.2, we forecast the volatility of day $t+1$. We also compare the result with a forecast made by fBm, using the same estimation sample.
@@
Figure 9 displays the trajectory of the realized volatility of the CAC 40 index, along with the dynamic estimates of the parameters $\alpha$ (between 1.65 and 2) and $H$ (lower than $1/2$) of an LFSM. We also represent $H-1/\alpha$, that is the memory of the process, which, as one can see in Figure 9, is always negative. It means that the sign of the forecast of the future increment of volatility is simply the opposite of the one of the last volatility increment, if the forecast is based on this sole lagged observation. In other words, when $d=2$ the LFSM and the fBm will lead to the same hit ratios.
@@
Beyond realized volatilities, we are also interested in the VIX series. We import from Bloomberg the daily close price of VIX between January 2000 and 12th April 2021. Empirical studies on rough volatility rarely consider this series of volatility. The Hurst exponent of the VIX series is indeed not similar to the one estimated for realized volatilities, with a value much closer to $1/2$ [74]. Figure 11 indeed shows that the time series of VIX looks like a smoothed version of the realized volatility of SPX. The Hurst exponent is between 0.1 and 0.6 and oscillates around 0.4. The parameter $\alpha$ is only slightly lower than 2. However, extreme events are not as rare as in a Gaussian distribution: the Jarque-Bera statistic above 10,000 and the kurtosis at 9.78 reject the Gaussian hypothesis with a confidence larger than $99.9\%$. There is also an asymmetry in the tails, with an EVT parameter on the right significantly above 0 (0.163) and an EVT parameter on the left not significantly different from 0 (-0.030).
Figure 11: Time series of VIX (divided by 100) in black and of the annualized realized volatility of the SPX index in grey (top left), estimated $\alpha$ (top right), $H$ (bottom left), and $H-1/\alpha$ (bottom right) of the VIX time series, using a two-year rolling window.
Since $H-1/\alpha$ is much closer to zero than in the rough volatility paradigm, forecasts are less accurate, consistently with the observations made in Section 4. Figure 12 shows that the fBm is outperforming when $d$ is 2 or 3, whereas the LFSM makes better predictions with $d$ equal to 6 or 8. Overall, the hit ratio is significantly above $1/2$ when $d$ is large enough, namely $d\geq 4$.
Figure 12: Hit ratio of the forecast of the next daily variation of VIX with an LFSM (black) or an fBm (grey), for $d\in\llbracket 2,12\rrbracket$. The two thin solid lines correspond to the confidence bound at $99\%$ and $99.9\%$ of a binomial test, for 4856 predictions.
@@ 5.2 High-frequency FX rates
Many papers have already studied the relevance of the fBm for modelling FX rates [10, 67], along with forecasts consistent with this model [24, 28]. Empirical literature also highlights the relevance of a stable distribution compared to a Gaussian model [26]. Figure 13 represents the three time series of FX rates. It clearly presents some jumps, which legitimize using a stable process instead of a Gaussian one. We then estimate the parameters of an LFSM on the log of each time series, using a rolling window of 720 observations, that is 12 hours. As one can see in Figure 13, the value of $\alpha$ is in general between $1.2$ and 2, with a peak below 1 for GBPUSD on the 26th June, $H$ fluctuates around $1/2$, with a negative peak of the estimator for GBPUSD on the 26th June. The value of the memory, $H-1/\alpha$, is more often negative than positive. We remark that H1/2H-1/2 and $H-1/\alpha$ do not always have the same sign. This indicates that the forecast using the LFSM is not always the same as the one using the fBm, even for $d=2$.